PR=2x, PQ=5x-4, and QR=x, determine the numerical length of PR
1
Expert's answer
2020-10-05T10:30:36-0400

We draw a horizonal line PR and point Q somewhere on the line.

"PQ=5x-4"

"QR=x"

"PR=2x"




"PQ+QR=PR"


"5x-4+x=2x"

Solve for x

"5x+x-2x=4"

"4x=4"

"x=1"

Now we substitute the value of x in PR.

"PR=2x"

"PR=2 \cdot 1=2"

Answer: 2.